An open box is to be constructed from a square piece of sheet metal by removing a square of side 3 feet from each corner and turning up the edges. If the box is to hold 300 cubic feet, what should be the dimensions of the sheet metal?
Answer by checkley77(12569)

Base=x^2
v=bh
300=3*x^2
3x^2=300
x^2=300/3
x^2=100
x=sqrt100
x=10 ft. for the sides of the base.
(10+2*3)=10+6=16 ft. is the sides of the original sheet.